# Minimal Thermodynamic Cost of Computing with Circuits

### v2

@@ V.2 MMC of a circuit family for heterogeneous priors
The homogeneous prior basis refers to a scenario where all logical gates in a circuit share the same prior distribution. However, this assumption is rarely valid in practice. The physical implementation of a logic gate often depends on its type; for instance, an AND gate typically has different underlying physical processes than an XOR gate. As a result, just like other physical properties of logic gates, the prior distribution naturally varies with gate type.
This heterogeneity has consequences for the thermodynamic cost of a circuit. Fig. 10 illustrates how mismatch cost changes when considering a circuit family composed of OR and AND gates. The figure demonstrates that the mismatch cost is consistently higher when the prior distribution of OR gates differs from that of AND gates across the circuit family.
@@
Fig. 12 illustrates how the mismatch cost of various individual logic gates changes as the prior distribution shifts. Notably, the mismatch cost increases as the prior distribution deviates further from the uniform distribution. Similarly, Fig. 13 shows how the total MMC of a circuit composed of AND and OR gates varies as the prior distributions of the two gate types become increasingly heterogeneous.
@@
Figure 12: Mismatch cost of individual gates as a function of their prior distributions. The input distribution to the gate is i.i.d uniform.
V.3 Smallest MMC and smallest size circuit implementing a given function
@@ VI Discussion and Future work
A key result of our study is Th. 1, which establishes an upper bound on the mismatch cost complexity that scales linearly with the size complexity of a circuit family under the assumption of a homogeneous prior distribution across all gates. Understanding precisely what causes the slope differences between circuit families, is still a question. In contrast to Th. 1, Th. 2 relaxes this assumption, allowing for cases where mismatch cost complexity deviates from size complexity. This result opens avenues for further exploration into how circuit structure, gate heterogeneity, and prior distributions influence thermodynamic costs. For example, with known prior distributions for each type of gate, a key question arises: how can we optimize the circuit topology associated with a Boolean function to minimize the thermodynamic cost?
@@
Another open question concerns the role of fan-in and fan-out—the number of inputs to, and outputs from, a gate, respectively—in determining the total mismatch cost. Fan-out greater than one allows the same gate output to be reused in multiple locations, often reducing the overall circuit size. However, it remains unclear how such reuse affects the minimal thermodynamic cost. Does increasing fan-out reduce the mismatch cost, just as it reduces size? Or could it increase or leave the cost unchanged? Investigating this trade-off between logical reuse and energetic dissipation is an important direction for future work.
@@
A fundamental problem is identifying the circuit family for a given Boolean function that minimizes mismatch cost. Understanding how this circuit family differs from those optimized for size or depth complexity would clarify the fundamental trade-offs between space, time, and thermodynamic cost in computation. Another intriguing direction is the thermodynamic analysis of circuit classes such as NC, which consists of functions that can be efficiently parallelized. A key open question is whether functions in NC—those that can be computed efficiently in parallel—are also implementable by circuits with low thermodynamic cost. Investigating this could provide new insights into the interplay between parallelizability and energetic efficiency in computation.
@@ VII Acknowledgment
This work was supported by the U.S. National Science Foundation (NSF) Grant 2221345. We thank the Santa Fe Institute for helping to support this research.
Appendix A Derivation of Eq. 24
@@ A.1 Subsystem process
If $q^{0}(x_{A})$ and $q^{0}(x_{B})$ are the prior distributions for the evolution of subsystems $A$ and $B$ respectively, the prior distribution for the joint evolution of A and B, such that 39 and 40 hold, is the product distribution,
@@ A.2 MMC lower bound on EP of a subsystem process
Consider a system $S$ consisting of multiple subsystems that evolve one after another, and the dynamics of one or more of these subsystems does not directly depend on the state of some other subsystems. In particular, let us focus on a subsystem $\mu$ that evolves based on the value of another subsystem denoted as $\mathrm{pa}(\mu)$. The set of the rest of the subsystems is denoted as $\tilde{\mu}$. This results in a particular partition of the set of subsystems, $S=\mu\cup\mathrm{pa}(\mu)\cup\tilde{\mu}$. Let us denote $p^{0}$ the joint probability distribution of the system before subsystem $\mu$ updates. Accordingly, let $p^{1}$ denote the joint probability distribution after subsystem $\mu$ updates, while the rest of the system remains the same. The distribution undergoes a change,
$p^{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)},\boldsymbol{x}_{\tilde{\mu}})\longrightarrow p^{1}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)},\boldsymbol{x}_{\tilde{\mu}}),$ (42)
such that $p^{1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\sum_{\boldsymbol{x}_{\tilde{\mu}}}p^{1}(\boldsymbol{x})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})p^{0}_{\mu}(\boldsymbol{x}_{\mu})$, and since the rest of the subsystems do not change, we have $p^{1}_{\tilde{\mu}}(\boldsymbol{x}_{\tilde{\mu}})=\sum_{\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)}}p^{1}(\boldsymbol{x})=p^{0}_{\tilde{\mu}}(\boldsymbol{x}_{\tilde{\mu}})$.
As mentioned above in A.1, the prior distribution for this kind of subsystem process is a product distribution of the form $q^{0}(\boldsymbol{x})=q^{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)}))q^{0}(\boldsymbol{x}_{\tilde{\mu}})$. Under the evolution of the subsystem $\mu$ conditioned on the state of subsystem $\mathrm{pa}(\mu)$, the prior distribution evolves to $q^{1}(\boldsymbol{x})=q^{1}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{0}(\boldsymbol{x}_{\tilde{\mu}})$.
$q^{0}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)}))q^{0}(\boldsymbol{x}_{\tilde{\mu}})\longrightarrow q^{1}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{0}(\boldsymbol{x}_{\tilde{\mu}})$ (43)
where $q^{1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{0}_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$.
@@
$\mathcal{MC}=D\left(p^{0}\|q^{0}\right)-D\left(p^{1}\|q^{1}\right)$ (44)
where the first KL divergence can be expanded to,
$D\left(p^{0}\|q^{0}\right)=\mathrm{I}_{0}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{0}_{\mu,\mathrm{pa}(\mu)}\|q^{0}_{\mu,\mathrm{pa}(\mu)}\right)+D\left(p^{0}_{\tilde{\mu}}\|q^{0}_{\tilde{\mu}}\right),$ (45)
@@
$D\left(p^{1}\|q^{1}\right)=\mathrm{I}_{1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{1}_{\mu,\mathrm{pa}(\mu)}\|q^{1}_{\mu,\mathrm{pa}(\mu)}\right)+D\left(p^{0}_{\tilde{\mu}}\|q^{0}_{\tilde{\mu}}\right).$ (46)
Since $p^{1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})p^{0}_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$ and $q^{1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{0}_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$, therefore, using chain rule for KL divergence, $D\left(p^{1}_{\mu,\mathrm{pa}(\mu)}\|q^{1}_{\mu,\mathrm{pa}(\mu)}\right)=D\left(p^{0}_{\mathrm{pa}(\mu)}\|q^{0}_{\mathrm{pa}(\mu)}\right)$. Therefore,
$D(p^{1}\|q^{1})=\mathrm{I}_{1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{0}_{\mathrm{pa}(\mu)}\|q^{0}_{\mathrm{pa}(\mu)}\right)+D\left(p^{0}_{\tilde{\mu}}\|q^{0}_{\tilde{\mu}}\right)$ (47)
@@
$\mathcal{MC}=\Delta\mathrm{I}+D\left(p^{0}_{\mu,\mathrm{pa}(\mu)}\|q^{0}_{\mu,\mathrm{pa}(\mu)}\right)-D\left(p^{0}_{\mathrm{pa}(\mu)}\|q^{0}_{\mathrm{pa}(\mu)}\right)$ (48)
where $\Delta\mathrm{I}=\mathrm{I}_{1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)-\mathrm{I}_{0}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)$ is the drop in mutual information between the subsystem consisting of $\boldsymbol{x}_{\mu}$ and $\boldsymbol{x}_{\mathrm{pa}(\mu)}$ and the rest of the system. Note that Eq. 48 provides an alternative derivation of the mutual information contribution to entropy production in systems with multiple communicating components—such as those involved in feedback control protocols—as previously derived in [26, 27] and [28]. Specifically, it is important to emphasizes that the contribution to entropy production arising from the making and breaking of correlations among components in a non-equilibrium process is captured by MMC.
@@ A.3 Application of subsystem MMC to circuits
When a gate or a layer of gates in a circuit is updated based on the state of the parent gates (or parent layers) while the rest of the circuit remains unchanged, it constitutes a subsystem process. Once again, let’s denote $\mu$ the gate or the layer of gates As described in Sec III.1.3, under the update of gate or layer $\mu$, the actual distribution of the circuit changes from $p^{\mu}(\boldsymbol{x})=p_{:\mu}(\boldsymbol{x}_{:\mu})p_{\mu,\mu:}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mu:})$ to $p^{\mu+1}(\boldsymbol{x})=p_{:\mu,\mu}(\boldsymbol{x}_{:\mu},\boldsymbol{x}_{\mu})p_{\mu:}(\boldsymbol{x}_{\mu:})$. The mutual information between the subsystem $\{\mu,\mathrm{pa}(\mu)\}$ and the rest of the circuit can be decomposed into the sum of two components: the mutual information between $\mu$ and its children $\mathrm{ch}(\mu)\subset\tilde{\mu}$, and the mutual information between $\mathrm{pa}(\mu)$ and its directly connected neighborhood of gates, which includes its parent $\mathrm{pa}(\mathrm{pa}(\mu))$ and its children $\mathrm{ch}(\mathrm{pa}(\mu))$,
$\mathrm{I}_{\mu}(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}})=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+\mathrm{I}_{\mu}(\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\mathrm{pa}(\mathrm{pa}(\mu))},\mathrm{X}_{\mathrm{ch}(\mathrm{pa}(\mu))})$ (49)
In the distribution $p^{\mu+1}(\boldsymbol{x})$, the variables $\mathrm{X}_{\mu}$ and $\mathrm{X}_{\mathrm{ch}(\mu)}$ are independent, implying that $\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})=0$. On the other hand, the mutual information between $\mathrm{pa}(\mu)$ and its directly connected neighborhood of gates remains the same in the distribution $p^{\mu+1}$ as in $p^{\mu}$. Therefore,
$\Delta\mathrm{I}$ $=\mathrm{I}_{\mu}(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}})-\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}})$ (50)
$=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)}).$ (51)
Using Eq. 48, the mismatch cost of updating gate $\mu$ is,
$\mathcal{MC}_{\mu}=\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)})\right)-D\left(p_{\mathrm{pa}(\mu)})||q_{\mathrm{pa}(\mu)}\right)$ (52)
The joint distribution of the circuit in a layer-by-layer implementation evolves in a manner similar to that of the gate-by-gate implementation (Fig. 4). The distribution of the circuit before the update of layer $l$ is,
$p^{l}(\boldsymbol{x})=p_{:l}(\boldsymbol{x}_{:l})\,p_{l,l:}(\boldsymbol{x}_{l},\boldsymbol{x}_{l:}).$ (53)
After the update of layer $l$, the joint distribution evolves to,
$p^{l+1}(\boldsymbol{x})=p_{:l,l}(\boldsymbol{x}_{:l},\boldsymbol{x}_{l})p_{l:}(\boldsymbol{x}_{l:}).$ (54)
The prior associated with the update of layer $l$ is defined similar to Eq. (21),
$q^{l}(\boldsymbol{x})=q_{l,\mathrm{pa}(l)}(\boldsymbol{x}_{l},\boldsymbol{x}_{\mathrm{pa}(l)})q_{-1}(\boldsymbol{x}_{-(l\cup\mathrm{pa}(l))})$ (55)
which after the update of layer $l$ evolves to,
$\tilde{q}^{l}(\boldsymbol{x})=\tilde{q}_{l,\mathrm{pa}(l)}(\boldsymbol{x}_{l},\boldsymbol{x}_{\mathrm{pa}(l)})q_{/(l\cup\mathrm{pa}(l))}(\boldsymbol{x}_{/(l\cup\mathrm{pa}(l))})$ (56)
The associated mismatch cost is,
$\mathcal{MC}_{l}$ $=D(p^{l}||q^{l})-D(p^{l+1}||\tilde{q}^{l})$ (57)
$=\mathrm{I}_{l}(\mathrm{X}_{l};\mathrm{X}_{\mathrm{ch}(l)})+D\left(p_{l}p_{\mathrm{pa}(l)}||q_{l,\mathrm{pa}(l)}\right)-D\left(p_{\mathrm{pa}(l)}||q_{\mathrm{pa}(l)}\right)$

### v3

@@ V.2 Relationship between MMC complexity and size complexity
Moreover, the assumption of a homogeneous prior is rarely realistic in practice. The physical implementation of a logic gate typically depends on its type; for example, an AND gate may involve different underlying physical processes than an XOR gate. Consequently, just like other physical characteristics of gates, the prior distribution naturally varies with gate type. This heterogeneity influences the thermodynamic cost of a circuit, as discussed above and illustrated in Fig. 9.
@@
Fig. 12 demonstrates how the mismatch cost of individual logic gates changes as their prior distributions shift. In particular, the mismatch cost tends to increase as the prior deviates further from uniform. Similarly, Fig. 13 shows the impact of heterogeneous priors on the total MMC of a fixed circuit composed of AND and OR gates, highlighting how increasing disparity between gate priors can affect the overall thermodynamic cost.
@@
Figure 12: Mismatch cost of individual gates as a function of their prior distributions. The input distribution to the gate is uniform. The x-axis represents the support of prior distribution at $x_{g}=1$.
V.3 Minimizing MMC or Minimizing circuit size for a given Boolean function?
@@ VI Discussion and Future work
Our central technical result for Boolean circuits is Eq. (25), which provides a closed-form expression for the total MMC incurred when a circuit is run on a given input distribution. This expression applies equally to gate-by-gate execution and to layer-by-layer execution. Once the prior distributions associated with the gates are specified—capturing, at a coarse-grained and design-level assumptions about typical operating conditions—the total MMC becomes a circuit-level quantity that can be evaluated directly from the circuit’s architecture. In this sense, MMC is an unavoidable contribution to dissipation that is rooted fundamentally in the topology and modular organization of the circuit itself, together with the mismatch between the input statistics encountered in practice and those for which the circuit is effectively optimized.
@@
Developed from Eq. 25, a key result of our study is Th. (1), which establishes an upper bound on the mismatch cost complexity that scales linearly with the size complexity of a circuit family under the assumption of a homogeneous prior distribution across all gates. Understanding precisely what causes the slope differences between circuit families, is still a question. In contrast to Th. (1), Th. (2) relaxes this assumption, allowing for cases where mismatch cost complexity deviates from size complexity. This result opens avenues for further exploration into how circuit structure, gate heterogeneity, and prior distributions influence thermodynamic costs. For example, with known prior distributions for each type of gate, a key question arises: how can we optimize the circuit topology associated with a Boolean function to minimize the thermodynamic cost?
Another key concept we introduce is mismatch cost complexity, which puts energetic cost on similar footing with more familiar resources like circuit size and depth. This lets us define corresponding mismatch cost complexity classes, which classifies computational problems according to whether they admit circuit families whose worst-case MMC grows no faster than some specified function. The motivation is very much in the spirit of standard complexity classes such as $\mathbf{P_{/poly}}$ or $NC$ (Nick’s class), except that here the organizing principle is thermodynamic cost rather than purely combinatorial resources. This viewpoint opens up a new direction of thinking about computation that is grounded in physical resource constraints. For instance, we showed that $\mathbf{P_{/poly}}\subseteq\mathbf{MMC}(\mathrm{poly})$: any problem solvable by polynomial-size circuits can also be implemented by circuits whose worst-case mismatch cost grows at most polynomially. What we do not yet know is whether the converse holds—namely, whether polynomially bounded MMC necessarily implies polynomial size, i.e., whether $\mathbf{P_{/poly}}=\mathbf{MMC}(\mathrm{poly})$. More broadly, the properties of mismatch cost complexity is still largely unexplored, and getting a better handle on it may shed new light on the energetic resource requirements of computation, beyond what size and depth alone can capture.
@@
Another open question concerns the role of fan-in and fan-out—the number of inputs to, and outputs from, a gate, respectively—in determining the total mismatch cost. Fan-out greater than one allows the same gate output to be reused at multiple downstream locations, often reducing overall circuit size. However, it remains unclear how such logical reuse affects the minimal thermodynamic cost. Does increasing fan-out tend to reduce MMC, in analogy with its effect on size? Or might it leave MMC unchanged—or even increase it—by inducing additional correlations loss with each run? Understanding this trade-off between logical reuse and energetic dissipation is not an important theoretical question but it has direct implications for circuit architecture and hardware design. More generally, our observation that layer-by-layer execution yields a lower MMC than gate-by-gate execution suggests a practical design principle: to the extent possible, implementing gates simultaneously in synchronized, clocked layers may reduce thermodynamic overhead relative to more sequential execution. This points to a potential energetic advantage of highly parallel, clocked architectures, beyond their well-known benefits for latency and throughput.
@@
Another fundamental problem is identifying the circuit family for a given Boolean function that minimizes mismatch cost. Understanding how this circuit family differs from those optimized for size or depth complexity would clarify the fundamental trade-offs between space, time, and thermodynamic cost in computation. Another intriguing direction is the thermodynamic analysis of circuit classes such as NC, which consists of functions that can be efficiently parallelized. A key open question is whether functions in NC—those that can be computed efficiently in parallel—are also implementable by circuits with low thermodynamic cost. Investigating this could provide new insights into the interplay between parallelizability and energetic efficiency in computation.
@@ VII Acknowledgment
This work was supported by the U.S. National Science Foundation (NSF) Grant 2221345. We thank the Santa Fe Institute for helping to support this research. We also thank Sosuke Ito, David Soloveichik and David Doty for helpful discussions.
Appendix A Derivation of Eq. (25)
@@ A.1 Subsystem process
If $q^{0}(x_{A})$ and $q^{0}(x_{B})$ are the prior distributions for the evolution of subsystems $A$ and $B$ respectively, the prior distribution for the joint evolution of A and B, such that condition (38) and (39) hold, is the product distribution,
@@ A.2 MMC lower bound on EP of a subsystem process
Consider a system $S$ consisting of multiple subsystems that evolve one after another, and the dynamics of one or more of these subsystems does not directly depend on the state of some other subsystems. In particular, let us focus on a subsystem $\mu$ that evolves based on the value of another subsystem denoted as $\mathrm{pa}(\mu)$. The set of the rest of the subsystems is denoted as $\tilde{\mu}$. This results in a particular partition of the set of subsystems, $S=\mu\cup\mathrm{pa}(\mu)\cup\tilde{\mu}$. Let $p$ and $\hat{p}$ denote the probability distribution of the joint system before and after the update of subsystem $\mu$, respectively, while the rest of the system remains unchanged. The distribution undergoes a change,
[…]
where $q^{\mu+1}_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=\pi_{\mu}(\boldsymbol{x}_{\mu}|\boldsymbol{x}_{\mathrm{pa}(\mu)})q^{\mu}_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$.
@@
$\mathcal{MC}_{\mu}=D\left(p^{\mu}\|q^{\mu}\right)-D\left(p^{\mu+1}\|q^{\mu+1}\right)$ (43)
where the first KL divergence can be written as,
$D\left(p^{\mu}\|q^{\mu}\right)=\mathrm{I}_{\mu}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{\mu}_{\mu,\mathrm{pa}(\mu)}\|q^{\mu}_{\mu,\mathrm{pa}(\mu)}\right)+D\left(p^{\mu}_{\tilde{\mu}}\|q^{\mu}_{\tilde{\mu}}\right),$ (44)
@@
$D\left(p^{\mu+1}\|q^{\mu+1}\right)=\mathrm{I}_{\mu+1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{\mu+1}_{\mu,\mathrm{pa}(\mu)}\|q^{\mu+1}_{\mu,\mathrm{pa}(\mu)}\right)+D\left(p^{\mu}_{\tilde{\mu}}\|q^{\mu}_{\tilde{\mu}}\right).$ (45)
[…]
$D(p^{\mu+1}\|q^{\mu+1})=\mathrm{I}_{\mu+1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p^{\mu}_{\mathrm{pa}(\mu)}\|q^{\mu}_{\mathrm{pa}(\mu)}\right)+D\left(p^{\mu}_{\tilde{\mu}}\|q^{\mu}_{\tilde{\mu}}\right)$ (46)
@@
[…]
As an aside, Eq. (47) provides an alternative derivation of the mutual information contribution to entropy production in systems with multiple communicating components—such as those involved in feedback control protocols—as previously derived in [40, 41] and [34]. Specifically, it emphasizes that the contribution to entropy production arising from the making and breaking of correlations among components in a non-equilibrium process is captured by the MMC.
Note: for simplicity we drop the superscripts (which indicates what step of the computation are we in) and we leave the subscripts (which indicates which gate (or layer) we are updating) since we update sequentially in (48).
@@ A.3 Application of subsystem MMC to circuits
When a gate or a layer of gates in a circuit is updated based on the state of the parent gates (or parent layers) while the rest of the circuit remains unchanged, it constitutes a subsystem process. Once again, let us use $\mu$ to denote a non-input gate or layer of gates. As described in Sec III.1.3, under the update of gate or layer $\mu$, the actual distribution of the circuit changes from $p^{\mu}(\boldsymbol{x})=p_{:\mu}(\boldsymbol{x}_{:\mu})p_{\mu,\mu:}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mu:})$ to $p^{\mu+1}(\boldsymbol{x})=p_{:\mu,\mu}(\boldsymbol{x}_{:\mu},\boldsymbol{x}_{\mu})p_{\mu:}(\boldsymbol{x}_{\mu:})$.
Using Eq. (47), the mismatch cost of running gate $\mu$ is given by the expression,
$\mathcal{MC}_{\mu}=\mathrm{I}_{\mu}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)-\mathrm{I}_{\mu+1}\left(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}}\right)+D\left(p_{\mu,\mathrm{pa}(\mu)}\|q_{\mu,\mathrm{pa}(\mu)}\right)-D\left(p_{\mathrm{pa}(\mu)}\|q_{\mathrm{pa}(\mu)}\right)$ (48)
[…]
$=S_{\mu}(\mathrm{X}_{\mu})+S_{\mu}(\mathrm{X}_{\mathrm{pa}(\mu)})+S_{\mu}(\mathrm{X}_{\tilde{\mu}})-S_{\mu}(\mathrm{X}),$ (51)
[…]
$\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)};\mathrm{X}_{\tilde{\mu}})=S_{\mu+1}(\mathrm{X}_{\mu},\mathrm{X}_{\mathrm{pa}(\mu)})+S_{\mu+1}(\mathrm{X}_{\tilde{\mu}})-S_{\mu+1}(\mathrm{X}).$ (52)
[…]
$\Delta\mathrm{I}_{\mu}$ $=\mathrm{I}_{\mu}-\mathrm{I}_{\mu+1}$ (53)
$=\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+S_{\mu+1}(\mathrm{X})-S_{\mu}(\mathrm{X}).$ (54)
When $\Delta\mathrm{I}_{\mu}$ contributions are summed over all the non-input gates in the circuit (in any topological order), we get:
$\sum_{\mu\in V}\Delta\mathrm{I}_{\mu}$ $=\sum_{\mu\in V_{\mathrm{nin}}}\left[\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+S_{\mu+1}(\mathrm{X})-S_{\mu}(\mathrm{X})\right]$ (55)
$=S_{|V_{\mathrm{nin}}|+1}(\mathrm{X})-S_{1}(\mathrm{X})+\sum_{\mu\in V_{\mathrm{nin}}}\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})$ (56)
Eq. (56) is consistent with earlier results on the loss of mutual information in Bayesian networks considered in [34]. [34] considers thermodynamic analysis of more general Bayesian networks with branches, loops, recurrence, etc. Note that $S_{1}(\mathrm{X})$ denote the Shannon entropy of the joint circuit state before any non-input gates are run on the new inputs, and $S_{|V_{\mathrm{nin}}|+1}(\mathrm{X})$ is the entropy after all non-input gates have completed their updates.
Before any non-input gates are executed, the values stored in the non-input nodes $\mathrm{X}_{\mathrm{nin}}$ are remnants of the previous run, while the input nodes $\mathrm{X}_{\mathrm{in}}$ are freshly sampled from the new input distribution. By construction of the model, these two collections of variables are initially statistically independent. Therefore,
$S_{1}(\mathrm{X})=S_{1}(\mathrm{X}_{\mathrm{nin}},\mathrm{X}_{\mathrm{in}})=S_{1}(\mathrm{X}_{\mathrm{nin}})+S_{1}(\mathrm{X}_{\mathrm{in}}).$ (57)
After the circuit has finished updating, the non-input nodes generally become correlated with the inputs through the computation. Hence, the final joint entropy can be written as
$S_{|V_{\mathrm{nin}}|+1}(\mathrm{X})=S(\mathrm{X}_{\mathrm{nin}},\mathrm{X}_{\mathrm{in}}),$ (58)
Combining these expressions, we obtain
$S_{1}(\mathrm{X})-S_{|V_{\mathrm{nin}}|+1}(\mathrm{X})=S_{1}(\mathrm{X}_{\mathrm{nin}})+S_{1}(\mathrm{X}_{\mathrm{in}})-S(\mathrm{X}_{\mathrm{nin}},\mathrm{X}_{\mathrm{in}})=\mathrm{I}_{0}(\mathrm{X}_{\mathrm{nin}};\mathrm{X}_{\mathrm{in}}).$ (59)
where $\mathrm{I}_{0}(\mathrm{X}_{\mathrm{nin}};\mathrm{X}_{\mathrm{in}})$ denotes the mutual information between the non-input and input nodes induced by the computation, before inputs are provided with new values and the entire circuit is correlated from the previous run. Therefore, the aggregate mismatch cost of running all non-input gates sequential is
$\sum_{\mu}\mathcal{MC}_{\mu}=-\mathrm{I}_{0}(\mathrm{X}_{\mathrm{nin}};\mathrm{X}_{\mathrm{in}})+\sum_{\mu\in V_{\mathrm{nin}}}\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+\sum_{\mu\in V_{\mathrm{nin}}}\left[D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)}\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right)\right]$ (60)
This mismatch cost of overwriting the input nodes with new values sampled from $p_{\mathrm{in}}$ is given by Eq. 19:
$\mathcal{MC}_{ow}(p_{\mathrm{in}})=\mathrm{I}_{0}(\mathrm{X}_{\mathrm{in}};\mathrm{X}_{\mathrm{nin}})+D(p_{\mathrm{in}}||q_{\mathrm{in}}).$ (61)
Total mismatch cost of the circuit is the sum of mismatch cost of overwriting the input nodes with new values and the aggregate mismatch cost of running all non-input gates. Thus, for an input distribution $p_{\mathrm{in}}$ and circuit $C_{n}$, the total mismatch cost is
$\mathcal{MC}(C_{n},p_{\mathrm{in}})$ $=\mathcal{MC}_{ow}(p_{\mathrm{in}})+\sum_{\mu\in V_{\mathrm{nin}}}\mathcal{MC}_{\mu}$ (62)
$=D(p_{\mathrm{in}}||q_{\mathrm{in}})+\sum_{\mu\in V_{\mathrm{nin}}}\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})+\sum_{\mu\in V_{\mathrm{nin}}}\left[D\left(p_{\mu}p_{\mathrm{pa}(\mu)}||q_{\mu,\mathrm{pa}(\mu)}\right)-D\left(p_{\mathrm{pa}(\mu)}||q_{\mathrm{pa}(\mu)}\right)\right].$ (63)